OpenStudy (mony01):

find the derivative of g(x)=2x+1/3x-1

OpenStudy (mathmale):

For the sake of clarity, please use parentheses in expressions such as 1/3x. It's hard to know whether you mean (1/3)*x or 1/(3x). Assuming that you mean (1/3)*x, the first power of x, with a constant, fractional coefficient, just differentiate each of the three terms of your expression separately: g'(x) = 2 + (1/3) = 7/3. Assuming that you mean 1/(3x), the reciprocal of 3x, rewrite g(x) as g(x) = 2x + (3x)^(-1) - 1 first. Apply the power and constant coefficient rules to each term as appropriate. Result: g'(x) = 2 -(3x)^(-2) (considerably different from before, no?)

OpenStudy (mony01):

oh ok i meant..\[g(x)=\frac{ 2x+1 }{ 3x-1 }\]

OpenStudy (mathmale):

Thanks. Your 2nd expression is so much clearer it's like night and day. Are you OK with finding the derivative of g(x) (using the quotient rule)?

OpenStudy (mony01):

pretty much i just am having trouble figuring out why the answer has a negative sign?

OpenStudy (mathmale):

I agree that the derivative of g(x) has a negative sign. Have you written out all of the steps of the quotient rule? (3x-1)(2) - (2x+1)(3) g'(x) = ------------------ (3x-1)^2 Try simplifying/evaluating this and see if you don't get a negative result also.

OpenStudy (mony01):

oh i did get a negative now, my bad i was doing (2x+1)(3)-(3x-1)(2)

OpenStudy (mathmale):

So you're OK now? Further questions?